the hight of a triangle is twice the base. the area is 49 square inches. find the base.
Answer by Pieter(35)

the formulae for the area of a triangle: (b*h)/2 with b=base and h=height
In your excercise we have: height = 2 * base or h = 2 * b
Putting this in the area-formulae: (b * 2 * b)/2 = (2*bē)/2 = bē
And this has to be 49: bē= 49
or b = sqrt(49) = 7
greets